OpenStudy (jdaisyflowerluv):

Solve for x: −3|2x + 6| = −12 A. x = 1 and x = 5 B. b. x = −1 and x = −5 C. x = −9 and x = 3 D.No solution I chose B but I am not sure

OpenStudy (irishboy123):

start with \(−3|2x + 6| = −12\) \(3|2x + 6| = 12\)

OpenStudy (loser66):

If you are not sure, just check one by one. Like this if the answer is A) then \(-3|2*1+6| ~~MUST ~= -12\) but the left hand side is -21 \(\neq 12\), then A is not the answer. Do the same with b) and c) If none of them is correct, then d) is the best choice. NOTE: I don't say that d is the answer.

OpenStudy (mathmale):

@JDaisyFlowerluv: Would you please share the work you have done on your own. When you have solved this problem, please close it (click on "Close," above). Thank you.

OpenStudy (ppery329):

The solution is B. You make two separate equations with the absolute value bars and make one of the solutions negative. Next, solve for x in each equation. You get -1 & -5.